# Robust a posteriori estimation of probit-lognormal seismic fragility curves via sequential design of experiments and constrained reference prior

### v2

@@ Abstract
A seismic fragility curve expresses the probability of failure of a structure conditional to an intensity measure (IM) derived from seismic signals. When only limited data is available, the practitioner often refers to the probit-lognormal model coupled with maximum likelihood estimation (MLE) to obtain estimates of these curves. This means that only a binary indicator of the state (BIS) of the structure is known, namely a failure or non-failure state indicator, when it is subjected to a seismic signal with an intensity measure IM. In this context, the objective of this work is to propose a method for optimally estimating such curves by obtaining the most precise estimate possible with the minimum of data. The novelty of our work is twofold. First, we present and show how to mitigate the likelihood degeneracy problem which is ubiquitous with small data sets and hampers frequentist approaches such as MLE. Second, we propose a novel strategy for sequential design of experiments (DoE) that selects seismic signals from a large database of synthetic or real signals via their IM values, to be applied to structures to evaluate the corresponding BISs. This strategy relies on a criterion based on information theory in a Bayesian framework. It therefore aims to sequentially designate the IM value such that the pair (IM, BIS) has on average, with respect to the BIS of the structure, the greatest impact on the posterior distribution of the fragility curve. The methodology is applied to a case study from the nuclear industry. The results demonstrate its ability to efficiently and robustly estimate the fragility curve, and to avoid degeneracy even with a limited amount of data. Furthermore, we demonstrate that the estimates quickly reach the model bias induced by the probit-lognormal modeling. Eventually, two criteria are suggested to help the user stop the DoE algorithm.
@@ 1 Introduction
In Section 2, the contributions and the positioning of this paper are first presented. The Bayesian framework is the subject of Section 3. In particular, the definitions of “likelihood degeneracy” and “robust a posteriori estimation” are given there. Section 4 is devoted to the presentation of our DoE strategy. Then, in Section 5 a comprehensive study of the performance of this strategy is presented considering a case study from the nuclear industry. The conclusion is finally proposed in Section 6. We specify that the results presented in the paper are also supported by the study of a “toy case” which is presented in A.
2 Contributions and positioning of this paper
This paper focuses on the estimation of seismic fragility curves using the lognormal model in the case where the data is sparse and binary, i.e., they only reflect the state of the structure (failure or non-failure) when subjected to seismic excitation. We base our data selection methodology on the use of a seismic signal generator (SSG) that allows us to enrich existing real signal databases corresponding to the seismic scenario of interest. Among the set of SSGs available in the literature, we used the parameterized stochastic model of modulated and filtered white-noise process defined in [56]. In our experience, this SSG efficiently addresses both temporal and spectral nonstationarities of seismic signals. From an engineering point of view, the current IMs of the generated artificial signals have statistical characteristics close to those of the real signals used for enrichment. This SSG has also been used in several works [24, 14, 57, 18, 22, 45, 15, 25].
@@ 2 Contributions and positioning of this paper
This article therefore mainly addresses the issues of equipment for which only binary seismic test results are available (for example, electrical relay qualification tests, etc.). Nevertheless, this also applies to simulation-based approaches whose results are reduced, by simple post-processing, to binary data. Numerical simulations provide much richer information than the simple binary state of a structure. Thus, they allow the use of more or less sophisticated statistical models, such as those mentioned in the introduction, to capture the relationship between the engineering demand parameter (EDP) and the IM, which can be complex, particularly when the behavior of the structure becomes non linear under the effect of earthquakes (which is the case of the industrial structure considered in this work). In such situations, when available data are limited, choosing the most appropriate statistical model is not a trivial task. In [21, 22, 23], the authors highlight that a major challenge for the application of surrogate models as part of earthquake engineering, is related to the random variability associated with seismic hazard which exhibits heteroscedastic behavior. This may result in using a large amount of data to obtain reliable estimates. In [14] and [15], the authors showed that the simplest model —thus involving minimal data— assuming a linear correlation between the logarithm of the EDP and that of the IM should be avoided due to the irreducible model bias that this can cause and which can greatly affect the estimation of fragility curves. In their setting, [14] showed that the probit-lognormal model gives better estimates than those based on linear regression. Although the use of a parametric model for the estimation of seismic fragility curves is subject to debate, its abundant use in the literature shows that the lognormal model appears to the practitioner as a model that is both pragmatic and relevant with a reduced number of data. Let us note that recent studies on the impact of IMs on fragility curves [18, 58, 59] suggest that the choice of an appropriate IM makes it possible to reduce the potential biases between reference fragility curves and their probit-lognormal estimations. Finally, let us add that regardless of the statistical model or strategy chosen to estimate seismic fragility curves, one way to minimize the amount of data to be used is to implement adaptive strategies. This is particularly true for structures with a high failure threshold relative to the seismic scenario of interest, even with the use of more sophisticated models than the probit-lognormal model [25].
@@
For all the reasons just mentioned, we seek in this paper to take full advantage of the probit-lognormal model by proposing a DoE strategy within a Bayesian framework. We take as support the reference prior theory in order to define an objective prior adapted to the probit-lognormal model. For that, we relied on the work of the reference [45] in which the authors derived the mathematical expression of the Jeffreys prior of the probit-lognormal model. Then, in order to facilitate the implementation and improve the numerical efficiency of the DoE algorithm, we propose a simplified analytical expression of the Jeffreys prior which takes up its main features. This expression also builds on recent developments presented in [60] that aim to slightly constrain the prior to ensure that it is proper. In our framework, this addresses the problems related to likelihood degeneracy, which are ubiquitous in small data sets and involve an improper posterior [45]. Since such posterior cannot be normalized as a probability, it prevents the use of MCMC algorithms. This is obviously a crucial point for any posterior estimation. These developments are presented in detail in this article and constitute, besides the DoE strategy itself, another major contribution of this work. Eventually, two criteria are suggested to help the user stop the DoE algorithm.
The proposed methodology is applied to a case study taken from the nuclear industry (a piping system). We use a numerical model that could today be described as a “low-fidelity model”, in order to have a large volume of data ($8\cdot 10^{4}$) allowing a comprehensive comparison of the performance of our strategy. Under our conditions, approximately 14% of the calculations involve nonlinear behavior of the piping system. This represents approximately 90 days of sequential calculations on a simple laptop. This numerical effort was carried out for comparative analysis purposes, but cannot be implemented in practice on the scale of an industrial installation. This is the objective of the proposed methodology, namely to enable a robust estimation of fragility curves involving a minimum of data.
@@
Finally, note that a new concept of robust fragility has emerged in the literature over the last two decades. It derives from the concept of robust reliability defined in [61] for the assessment of the reliability of structures. In this framework, “robustness” means that modeling uncertainties are explicitly taken into account, so that the calculated reliability is not sensitive to them. In the context of this paper, the robust fragility curve refers to the average fragility curve obtained by integration over the possible values of the parameters of the lognormal model, whose density is defined by the posterior distribution, and the associated credibility interval is quantified by its variance [62]. Knowing that with a limited amount of data, the probability of having a degenerate likelihood can be significant and that this can compromise any a posteriori estimation (the posterior can be improper, it depends on the prior) we propose to define a robust posterior estimation as an estimation based on a proper posterior. The definition that we propose presents the conditions that must be satisfied for the posterior to be proper. It therefore defines the conditions necessary for the estimation of a robust fragility curve.
@@ 3.2 Degeneracy of the likelihood
Figure 1: Example of a type 3 data sample $(\mathbf{z}^{k},\mathbf{a}^{k})$ which gives a degenerate likelihood. Graphs of (left) the likelihood given the degenerate data sample as a function of the tuple $(\alpha,\beta)$ and (right) the fragility curve (blue curve) according to which the points (red crosses) are sampled : each red cross is a tuple $(z,a)$, where its X-axis equals $a$ and its Y-axis equals $z$. On both figures, $a_{1}$ is the maximal observed IM value among “non-failures”, and $a_{2}$ is minimal one among “failures”.
@@ 3.3 Constrained reference prior and posterior
    with $\mathcal{F}(\theta)=-\sum_{z\in\{0,1\}}\int_{a\in\mathcal{A}}\ell_{1}(z,a|\theta)\nabla^{2}_{\theta}\log\ell_{1}(z,a|\theta)da.$
@@
      $\forall\alpha>0,\,\pi(\theta)\underset{\beta\rightarrow 0}{=}O(1),$ and $\pi(\theta)$ is integrable in the neighborhood of $\beta\longrightarrow\infty$.
@@
where $\mu_{A}$ and $\sigma_{A}$ respectively denote the mean and the standard deviation of the r.v. $\log A$. We let the reader notice that the distribution of the IM —which is approximated here by a lognormal distribution (see [45, 46] and Fig. 2)— is involved in the prior expression. As a matter of fact, its presence within the likelihood expression (Eq. (2)) propagates within the derivation of the Jeffreys prior, and so to its decay rates, as proven in [45]. The close connection that exists between the prior distribution of $\alpha$ and the one of $A$ emphasizes the need of taking into account a realistic distribution of the IM.
@@ 4.2 Approximation of the index
$=\int_{\Theta}\Psi^{z_{k+1}}_{a_{k+1}}(\theta)\prod_{q=p+1}^{k}\Psi^{z_{q}}_{a_{q}}(\theta)p(\theta|\mathbf{z}^{p},\mathbf{a}^{p})\frac{1}{L_{p}^{k}}d\theta,$
@@
$=\int_{\Theta}\Psi^{z_{k+1}}_{a_{k+1}}(\theta)^{1-\delta}\prod_{q=p+1}^{k}\Psi^{z_{q}}_{a_{q}}(\theta)p(\theta|\mathbf{z}^{p},\mathbf{a}^{p})\frac{1}{L^{k}_{p}}d\theta.$
@@
The above being proportional to the quantity $(\delta(\delta-1))^{-1}\left((Q^{0}_{1})^{\delta}Q^{0}_{1-\delta}+(Q^{1}_{1})^{\delta}Q^{1}_{1-\delta}\right)$. Eventually, $a_{k+1}$ is chosen as the maximal argument of the right-hand side term in the above equation.
@@ 4.3 Stopping criterion
Algorithm 1 consists in a loop where $k$ iterates from $k_{0}$ to $k_{\max}-1$. The upper value $k_{\max}$ represents the total number of experiments at the end of the campaign. In practical studies, that value is limited by the cost of experiments. As a matter of fact, additional experiments are expected to provide a quantity of information that enhances the estimates. Practitioners could judge to cease the campaign if the cost of an experiment overtakes its benefits. Such quantity of information provided by data samples is derived through our method in the index $\mathcal{I}$, that is why we suggest to study its variation to elucidate a stopping criterion:
@@ 5.1.1 Seismic signals generation
Figure 2: Histograms of IMs derived from $10^{5}$ generated synthetic signals : the IM is the PSA (left) and the PGA (right). They are compared with the densities of IMs coming from real accelerograms estimated by Gaussian kernel estimation (dashed lines), and with the densities of a lognormal distribution with same median and same log deviation (solid lines).
@@ 5.1.2 Description of the mechanical equipment
In this case study, we investigate the seismic fragility of a piping system that was tested on the Azalee shaking table at the EMSI laboratory of CEA/Saclay, as shown in Fig. 3-left. Fig. 3-right depicts the finite element model (FEM), based on beam elements and implemented through the proprietary FE code CAST3M [70]. The validation of the FEM was carried out thanks to an experimental campaign described in [71].
@@
Figure 3: (left) Overview of the piping system —which is part of a French pressurized water reactor— on the Azalee shaking table and (right) associated finite element model.
@@ 5.1.4 Sequential design of experiments implementation
5.2 Estimates and benchmarking metrics
Estimates of seismic fragility curves given observation $(\mathbf{z}^{k},\mathbf{a}^{k})$ are obtained by sampling independent and identically distributed values of $\theta$ from the posterior distribution $p(\theta|\mathbf{z}^{k},\mathbf{a}^{k})$. That sampling can be done via MCMC methods. In our work we use an adaptive Metropolis-Hastings algorithm [73] which necessitates iterative evaluations of the posterior up to a constant. For a given value of $\gamma$, we can evaluate the posterior from the approximation $\tilde{\pi}^{\ast}_{\gamma}$ suggested in Eq. (4). In [45], a more exact but more expensive computation of the Jeffreys prior is proposed. We precise that we have verified that the results obtained from their method to compute $\pi^{\ast}_{\gamma}$ or from the approximation $\tilde{\pi}^{\ast}_{\gamma}$ are indiscernible. Their comparison is not the point of this study and is not discussed in the following.

### v3

@@ Abstract
A seismic fragility curve expresses the probability of failure of a structure conditional to an intensity measure (IM) derived from seismic signals. When only limited data is available, the practitioner often refers to the probit-lognormal model coupled with maximum likelihood estimation (MLE) to obtain estimates of these curves. This means that only a binary indicator of the state (BIS) of the structure is known, namely a failure or non-failure state indicator, when it is subjected to a seismic signal with an intensity measure IM. In this context, the objective of this work is to propose a method for optimally estimating such curves by obtaining the most precise estimate possible with the minimum of data. The novelty of our work is twofold. First, we present and show how to mitigate the likelihood degeneracy problem which is ubiquitous with small data sets and hampers frequentist approaches such as MLE. Second, we propose a novel strategy for sequential design of experiments (DoE) that selects seismic signals from a large database of synthetic or real signals via their IM values, to be applied to structures to evaluate the corresponding BISs. This strategy relies on a criterion based on information theory in a Bayesian framework. It therefore aims to sequentially designate the IM value such that the pair (IM, BIS) has on average, with respect to the BIS of the structure, the greatest impact on the posterior distribution of the fragility curve. The methodology is applied to a case study from the nuclear industry. The results demonstrate its ability to efficiently and robustly estimate the fragility curve, and to avoid degeneracy even with a limited amount of data, i.e., less than 100. Furthermore, we demonstrate that the estimates quickly reach the model bias induced by the probit-lognormal modeling. Eventually, two criteria are suggested to help the user stop the DoE algorithm.
@@ 1 Introduction
In Section 2, the positioning, contribution and methodology of this paper are first presented. The Bayesian framework is the subject of Section 3. In particular, the definitions of “likelihood degeneracy” and “robust a posteriori estimation” are given there. Section 4 is devoted to the presentation of our DoE strategy. Then, in Section 5 a comprehensive study of the performance of this strategy is presented considering a case study from the nuclear industry. The conclusion is finally proposed in Section 6. We specify that the results presented in the paper are also supported by the study of a “toy case” which is presented in A.
2 Positioning, contribution and methodology
2.1 Positioning
@@
This paper focuses on the estimation of seismic fragility curves using the lognormal model in the case of limited (typically less than 100) and binary data, i.e., they only reflect the state of the structure (failure or non-failure) when subjected to seismic excitation. Therefore, it mainly addresses the issues of equipment for which only binary seismic test results are available (for example, electrical relay qualification tests, etc.). Nevertheless, this also applies to simulation-based approaches whose results are reduced, by simple post-processing, to binary data. Numerical simulations provide much richer information than the simple binary state of a structure. Thus, they allow the use of more or less sophisticated statistical models, such as those mentioned in the introduction, to capture the relationship between the engineering demand parameter (EDP) and the IM, which can be complex, particularly when the behavior of the structure becomes non linear under the effect of earthquakes (which is the case of the industrial structure considered in this work). In such situations, when available data is limited, choosing the most appropriate statistical model is not a trivial task. In [21, 22, 23], the authors highlight that a major challenge for the application of surrogate models as part of earthquake engineering, is related to the random variability associated with seismic hazard which exhibits heteroscedastic behavior. This may result in using a large amount of data to obtain reliable estimates. In [14] and [15], the authors showed that the simplest model —involving less data to adjust its parameters— assuming a linear correlation between the logarithm of the EDP and that of the IM should be avoided due to the irreducible model bias that it can introduce and which can greatly affect the estimation of fragility curves. In their setting, [14] showed that the probit-lognormal model gives better estimates than those based on linear regression. Although the use of a parametric model for the estimation of seismic fragility curves is subject to debate, its abundant use in the literature shows that the lognormal model appears to the practitioner as a model that is both pragmatic and relevant with a reduced number of data. Let us note that recent studies on the impact of IMs on fragility curves [18, 56, 57] suggest that the choice of an appropriate IM makes it possible to reduce the potential biases between reference fragility curves and their probit-lognormal estimations. Finally, let us add that regardless of the statistical model or strategy chosen to estimate seismic fragility curves, one way to minimize the amount of data to be used is to implement adaptive strategies. This is particularly true for structures with a high failure threshold relative to the seismic scenario of interest, even with the use of more sophisticated models than the probit-lognormal model [25].
2.2 Contribution
@@
For all the reasons just mentioned, we seek in this paper to take full advantage of the probit-lognormal model by proposing a DoE strategy within a Bayesian framework. We take as support the reference prior theory in order to define an objective prior adapted to the probit-lognormal model. For that, we relied on the work [45] in which the authors derived the mathematical expression of the Jeffreys prior of the probit-lognormal model. Then, in order to facilitate the implementation and improve the numerical efficiency of the DoE algorithm, we propose a simplified analytical expression of the Jeffreys prior which takes up its main features. This expression also builds on recent developments presented in [58] that aim to slightly constrain the prior to ensure that it is proper. In our framework, this addresses the problems related to likelihood degeneracy, which are ubiquitous in small data sets and involve an improper posterior [45]. Since such posterior cannot be normalized as a probability, it prevents the use of MCMC algorithms. This is obviously a crucial point for any posterior estimation. These developments are presented in detail in this article and constitute, besides the DoE strategy itself, another major contribution of this work. Eventually, two criteria are suggested to help the user stop the DoE algorithm.
@@
Note that a new concept of robust fragility has emerged in the literature over the last two decades. It derives from the concept of robust reliability defined in [59] for the assessment of the reliability of structures. In this framework, “robustness” means that modeling uncertainties are explicitly taken into account, so that the calculated reliability is not sensitive to them. In the context of this paper, the robust fragility curve refers to the average fragility curve obtained by integration over the possible values of the parameters of the lognormal model, whose density is defined by the posterior distribution, and the associated credibility interval is quantified by its variance [60]. Knowing that with a limited amount of data, the probability of having a degenerate likelihood can be significant and that this can compromise any a posteriori estimation (the posterior can be improper, it depends on the prior) we propose to define a robust posterior estimation as an estimation based on a proper posterior. The definition that we propose presents the conditions that must be satisfied for the posterior to be proper. It therefore defines the conditions necessary for the estimation of a robust fragility curve.
2.3 Methodology : principle and validation
Our methodology aims to optimally select seismic signals from a wide range of signals corresponding to a seismic scenario of interest. Since real signal databases are generally not rich enough to allow such a selection, we use in the first place a seismic signal generator (SSG) to enrich them. Among the set of SSGs available in the literature, we used the parameterized stochastic model of modulated and filtered white-noise process defined in [61]. This generator has been used several times since 2010 [24, 14, 62, 18, 22, 45, 15, 25]. In our experience, it efficiently addresses both temporal and spectral nonstationarities of seismic signals. From an engineering point of view, the current IMs of the generated artificial signals have statistical characteristics close to those of the real signals used for enrichment. For the present work, we rely on the $10^{5}$ synthetic signals generated by Sainct et al. [18]. They were obtained by calibrating the SSG from 97 real accelerograms selected in the European Strong Motion Database for a magnitude $M$ such that $5.5\leq M\leq 6.5$, and a source-to-site distance $R<20$ km [63].
It should be note that the DoE algorithm proposed in this work is independent of the generator and the seismic scenario. The study of the influence of the SSG is out of the scope of this work, which aims to propose a general methodological framework for the selection of seismic signals, for an optimal estimation of fragility curves.
For illustrative and performance evaluation purposes, the proposed methodology is applied to a case study taken from the nuclear industry: a piping system. We use here a numerical model based on beam elements, validated by comparison with tests performed on the Azalee shaking table at CEA/Saclay [64]. This model that could today be described as a "low-fidelity model", offers a good trade-off between accuracy and computation time. It allowed us to perform $8\cdot 10^{4}$ calculations, by randomly selecting $8\cdot 10^{4}$ excitation signals from among the $10^{5}$ artificial signals. Such a quantity of calculations would not have been reasonably achievable with a more sophisticated numerical model. In our setting, approximately 14% of these calculations involve nonlinear behavior (material nonlinearity) of the piping system. This represents approximately 90 days of sequential calculations on a simple laptop. This numerical effort was carried out for comparative analysis purposes, but cannot be implemented in practice on the scale of an industrial installation. That is, in fact, the objective of the proposed methodology, namely to enable a robust estimation of fragility curves involving a minimum of data (i.e., less than 100).
Thus, the $8\cdot 10^{4}$ calculations allow estimation of a reference fragility curve, against which the estimates of our method can be compared. This large number of results also allows the data selection methodology to be replicated many times, considering subsets of smaller size than the total validation sample. Each replication consists of randomly choosing two initial pairs (IM, BIS) and adding new pairs by sequentially applying the DoE algorithm up to 250 selections. Compared to a standard sampling method, these replications allow for statistical validation of the DoE method, based on performance metrics, such as the mean squared error, etc.
@@ 3.2 Degeneracy of the likelihood
Graphs of (left) the likelihood given the degenerate data sample as a function of the tuple $(\alpha,\beta)$ and (right) the fragility curve (blue curve) according to which the points (red crosses) are sampled : each red cross is a tuple $(z,a)$, where its X-axis equals $a$ and its Y-axis equals $z$. On both figures, $a_{1}$ is the maximal observed IM value among “non-failures”, and $a_{2}$ is the minimal one among “failures”.
Figure 1: Example of a type 3 data sample $(\mathbf{z}^{k},\mathbf{a}^{k})$ which gives a degenerate likelihood.
@@ 3.3 Constrained reference prior and posterior
    $\text{with}\quad\mathcal{F}(\theta)$ $=-\sum_{z\in\{0,1\}}\int_{a\in\mathcal{A}}\ell_{1}(z,a|\theta)\nabla^{2}_{\theta}\log\ell_{1}(z,a|\theta)da.$
@@
      $\forall\alpha>0,\,\pi(\theta)\underset{\beta\rightarrow 0}{=}0(1)$, and $\pi(\theta)$ is intgrable in the neighborhood of $\beta\rightarrow\infty$.
@@
where $\mu_{A}$ and $\sigma_{A}$ respectively denote the mean and the standard deviation of the r.v. $\log A$. We let the reader notice that the distribution of the IM —which is approximated here by a lognormal distribution— is involved in the prior expression. As a matter of fact, its presence within the likelihood expression (Eq. (2)) propagates within the derivation of the Jeffreys prior, and so to its decay rates, as proven in [45].
The close connection that exists between the prior distribution of $\alpha$ and the one of $A$ emphasizes the need of taking into account a realistic distribution of the IM. Approximating the IM distribution by a lognormal distribution is consistent with the distributions of the real and artificial signals of our study (see [45, 46] and Fig. 2). Some studies also show that a lognormal distribution is compatible with a scenario involving near-source ground motions [69]. In addition, this assumption is adopted for the development of ground motion prediction equations (GMPEs) [70, 71], from which artificial signals can be generated [72]. The scope of application of the prior defined in Eq. (4) is therefore quite broad.
@@ 4.2 Approximation of the index
$=\int_{\Theta}\Psi^{z_{k+1}}_{a_{k+1}}(\theta)\prod_{j=q+1}^{k}\Psi^{z_{j}}_{a_{j}}(\theta)p(\theta|\mathbf{z}^{q},\mathbf{a}^{q})\frac{1}{L_{q}^{k}}d\theta,$
@@
$=\int_{\Theta}\Psi^{z_{k+1}}_{a_{k+1}}(\theta)^{1-\delta}\prod_{j=q+1}^{k}\Psi^{z_{j}}_{a_{j}}(\theta)p(\theta|\mathbf{z}^{q},\mathbf{a}^{q})\frac{1}{L^{k}_{q}}d\theta.$
@@
The above being proportional to $(\delta(\delta-1))^{-1}\left((Q^{0}_{1})^{\delta}Q^{0}_{1-\delta}+(Q^{1}_{1})^{\delta}Q^{1}_{1-\delta}\right)$. Eventually, $a_{k+1}$ is chosen as the maximal argument of the right-hand side term in the above equation.
@@ 4.3 Stopping criterion
Algorithm 1 consists in a loop where $k$ iterates from $k_{0}$ to $k_{\max}-1$. The upper value $k_{\max}$ represents the total number of experiments at the end of the campaign. In practical studies, that value is limited by the cost of experiments. As a matter of fact, additional experiments are expected to provide a quantity of information that enhances the estimates. Practitioners could judge to cease the campaign if the cost of an experiment overtakes its benefits.
Such quantity of information provided by data samples is derived through our method in the index $\mathcal{I}$, that is why we suggest to study its variation to elucidate a stopping criterion:
@@ 5.1.1 Seismic signals generation
Histograms (in blue) of IMs (PSA in left figure and PGA in right figure) derived from $10^{5}$ generated synthetic signals. They are compared with the densities of IMs coming from real accelerograms estimated by Gaussian kernel estimation (dashed lines), and with the densities of lognormal distributions with same medians and same log deviations (solid lines).
Figure 2: Statistical distributions of the PSA and the PGA.
@@ 5.1.2 Description of the mechanical equipment
In this case study, we investigate the seismic fragility of a piping system that is part of a French pressurized water reactor and that was tested on the Azalee shaking table at the EMSI laboratory of CEA/Saclay, as shown in Fig. 3-left. Fig. 3-right depicts the finite element model (FEM), based on beam elements and implemented through the proprietary FE code CAST3M [75]. The validation of the FEM was carried out thanks to an experimental campaign described in [64].
@@
Figure 3: Overview of the piping system.
@@ 5.1.4 Sequential design of experiments implementation
5.2 Fragility curves estimations, benchmarking metrics and model bias
5.2.1 Fragility curves estimations
Estimations of seismic fragility curves given observation $(\mathbf{z}^{k},\mathbf{a}^{k})$ are obtained by sampling independent and identically distributed values of $\theta$ from the posterior distribution $p(\theta|\mathbf{z}^{k},\mathbf{a}^{k})$. That sampling can be done via MCMC methods. In our work we use an adaptive Metropolis-Hastings algorithm [77] which necessitates iterative evaluations of the posterior up to a constant. For a given value of $\gamma$, we can evaluate the posterior from the approximation $\tilde{\pi}^{\ast}_{\gamma}$ suggested in Eq. (4). In [45], a more exact but more expensive computation of the Jeffreys prior is proposed. We precise that we have verified that the results obtained from their method to compute $\pi^{\ast}_{\gamma}$ or from the approximation $\tilde{\pi}^{\ast}_{\gamma}$ are indiscernible. Their comparison is not the point of this study and is not discussed in the following.
5.2.2 Benchmarking metrics